You’re trying to determine whether or not to expand your business by building a new manufacturing plant. The plant has an installation cost of $10.8 million, which will be depreciated straight-line to zero over its four-year life. If the plant has projected net income of $1,293,000, $1,725,000, $1,548,000, and $1,310,000 over these four years, what is the project’s average accounting return (AAR)?
Answer #1

Average net income=Total net income/Total time period

=(1,293,000+1,725,000+1,548,000+1,310,000)/4=1469000

Average investment=(10,800,000/2)=$5,400,000

AAR=Average net income/Average investment

=1469000/5,400,000

=27.20%(Approx).